OpenStudy (anonymous):

sin25 cos35+cos25 sin35

jimthompson5910 (jim_thompson5910):

Hint: use the identity sin(x+y) = sin(x)cos(y) + cos(x)sin(y)

OpenStudy (anonymous):

how would I combine them?

OpenStudy (aum):

You can write the above identity in reverse: sin(x)cos(y) + cos(x)sin(y) = sin(x+y) Compare this to: sin25 cos35+cos25 sin35 Identify x and y. Plug the numbers on the right hand side of the identity.